Find the area bounded by the curve y=4x-x^2 and the lines x=-2 and y=4.


1
Expert's answer
2021-03-30T16:14:03-0400

Let us find the area bounded by the curve "y=4x-x^2" and the lines "x=-2" and "y=4". Firstly, let us sketch the graph:





It follows that that the area "A" is:

"A=\int_{-2}^2(4-(4x-x^2))dx=\int_{-2}^2(x^2-4x+4)dx=(\frac{x^3}{3}-2x^2+4x)|_{-2}^2=
\frac{8}{3}-8+8-(-\frac{8}{3}-8-8)=\frac{8}{3}+\frac{8}{3}+16=21\frac{1}{3}"


Answer: "21\frac{1}{3}" sq. units